### v1

@@ Discussion
Our results operate at the interface of the logic of belief and the physics of measurement in a quantum setting. We establish that the quantum world supports an analog of the Agreement Theorem of classical epistemics (Aumann, 1976) in the commuting case, while we find the distinctive epistemic phenomenon of CCD in the non-commuting regime. We therefore identify a clear boundary between classical intersubjectivity (ruled our under common certainty) and quantum intersubjectivity (allowed under common certainty). We show that CCD reduces to agreement under classical recording, offering a way to understand the classical Agreement Theorem in a quantum world.
@@
Importantly, QM bounds the disagreement that is possible across agents, via our $0\operatorname{-}1$ impossibility theorem. This result validates that our epistemic analysis has physical bite – there are epistemic states ruled out by QM. The impossibility result also distinguishes our approach from Frauchiger and Renner (2018). This latter work shows that chaining together statements of Alice’s certainty across mutually incompatible measurement branches yields a scenario that looks exactly like $0\operatorname{-}1$ disagreement, which is then – very reasonably – argued to suggest that conceptual inconsistencies may lurk in QM. By insisting that Alice’s certainties refer to the same underlying quantum state, we show, via our impossibility result, that the Frauchiger-Renner scenario cannot arise.
@@
We still allow disagreement in QM, albeit of a milder nature than $0\operatorname{-}1$ disagreement. So, we see our findings as challenging the notion that intersubjective disagreement should not be allowed in physics and, in this way, push in a similar direction to the QBist and relational accounts of QM (Fuchs and Schack, 2013; Rovelli, 1996). By introducing the tools of formal epistemics, our approach may suggest new treatments of these approaches.
@@
In previous work, Contreras-Tejada et al. (2021) formulate the agreement question in terms of a no-signaling box. Within this environment, a quantum agreement theorem follows and failed only for superquantum boxes. Liu, Chung, and Ramanathan (2024) generalize this result, while remaining in the commuting regime, to establish an agreement theorem for all quantum and almost-quantum regions. The present paper extends Contreras-Tejada et al. (2021) so that Alice’s and Bob’s operators can inhabit distinct, possibly non-commuting, operator algebras, and, as we have seen, this admits CCD even in a quantum setting. Leifer and Duarte (2022) treat the knowledge modality of the original classical Agreement Theorem, rather than the certainty modality, and establish the impossibility of common knowledge of disagreement in the setting of generalized probability theory (Barrett, 2007). (We comment on the difference between these modalities in the Supplementary Material.) Khrennikov and Basieva (2014) and Khrennikov (2015) study quantum-like agents interacting with quantum systems and are able to obtain CCD in this manner.
@@
Figure 2 shows that common certainty of probability assessments need not coincide with the shared assessment Alice and Bob would obtain after pooling their information. Here, Alice’s (resp. Bob’s) partition is the rows (resp. columns). The event of interest $E$ is $\{\omega_{1},\omega_{4}\}$. At any state, it is common certainty that Alice and Bob both assign probability $1/2$ to $E$. Yes, if they pooled their information, their shared assessment would be $1$ or 0, depending on the true state.
@@
Since $C_{*}=A_{*}B_{*}$ is the projector onto $\operatorname{ran}(A_{*})\cap\operatorname{ran}(B_{*})$, we have $0\leq C_{*}\leq A_{*}\leq A_{n}$ in the Loewner (positive semidefinite) order, and similarly $0\leq C_{*}\leq B_{*}\leq B_{n}$. Monotonicity of the trace with respect to the operator order and $\rho\geq 0$ give the inequalities and strict positivity. ∎
Lemma S.3 ensures that all conditional probabilities in the certainty recursion are well defined (no vanishing denominators).

### v2

@@ Discussion
Our results operate at the interface of the logic of belief and the physics of measurement in a quantum setting. We establish that the quantum world supports an analog of the Agreement Theorem of classical epistemics (Aumann, 1976) in the commuting case, while we find the distinctive epistemic phenomenon of CCD in the non-commuting regime. We therefore identify a clear boundary between classical intersubjectivity (ruled out under common certainty) and quantum intersubjectivity (allowed under common certainty). We show that CCD reduces to agreement under classical recording, offering a way to understand the classical Agreement Theorem in a quantum world.
@@
Importantly, QM bounds the disagreement that is possible across agents, via our $0\operatorname{-}1$ impossibility theorem. This result validates that our epistemic analysis has physical bite – there are epistemic states ruled out by QM. The impossibility result also distinguishes our approach from Frauchiger and Renner (2018). Their work shows that chaining together statements of Alice’s certainty across mutually incompatible measurement branches yields a scenario that looks exactly like $0\operatorname{-}1$ disagreement, which is then – very reasonably – argued to suggest that conceptual inconsistencies may lurk in QM. By insisting that Alice’s certainties refer to the same underlying quantum state, we show, via our impossibility result, that the Frauchiger-Renner scenario cannot arise.
@@
We still allow disagreement in QM, albeit of a milder nature than $0\operatorname{-}1$ disagreement. So, we see our findings as challenging the notion that intersubjective disagreement should not be allowed in physics and, in this way, push in a direction broadly similar to the QBist and relational accounts of QM (Fuchs and Schack, 2013; Rovelli, 1996). By introducing the tools of formal epistemics, our approach may suggest new treatments of these approaches.
@@
In previous work, Contreras-Tejada et al. (2021) formulate the agreement question in terms of a no-signaling box. Within this environment, a quantum agreement theorem follows and fails only for superquantum boxes. Liu, Chung, and Ramanathan (2024) generalize this result, while remaining in the commuting regime, to establish an agreement theorem for all quantum and almost-quantum regions. The present paper extends Contreras-Tejada et al. (2021) so that Alice’s and Bob’s operators can inhabit distinct, possibly non-commuting, operator algebras, and, as we have seen, this admits CCD even in a quantum setting. Leifer and Duarte (2022) treat the knowledge modality of the original classical Agreement Theorem, rather than the certainty modality, and establish the impossibility of common knowledge of disagreement in the setting of generalized probability theory (Barrett, 2007). (We comment on the difference between these modalities in the Supplementary Material.) Khrennikov and Basieva (2014) and Khrennikov (2015) study quantum-like agents interacting with quantum systems and are able to obtain CCD in this manner.
A natural extension of our framework would enrich the present protocol-agnostic model by a classical variable encoding the order of measurement by Alice and Bob, together with a description of the agents’ beliefs about that order, beliefs about beliefs about that order, and so on to higher levels as required. In Example 2, such an extension is immediate because Alice’s and Bob’s local measurements commute and order affects only the state Bob uses to form his probability estimate of $E$. If Bob believes that Alice has already performed her measurement, his relevant state is:
$\bar{\rho}_{A}=\sum_{i}P_{A}^{i}\,\rho\,P_{A}^{i},$ (37)
from which he estimates the probability of $E$ to be $1/2$ not $1$. Letting $\pi$ denote the probability, given outcome $j=0$, that Bob assigns to Alice’s having already measured, we obtain:
$q_{B}=\pi\cdot\dfrac{1}{2}+(1-\pi)\cdot 1=1-\dfrac{\pi}{2}.$ (38)
Bob’s posterior interpolates between the protocol-agnostic value $1$ we calculated earlier and the order-dependent value $1/2$. In a fuller epistemic model, Alice may be uncertain about Bob’s parameter $\pi$, Bob may be uncertain about Alice’s beliefs about $\pi$, and so on. Summing up, the quantum model supplies the order-dependent posteriors, while a conventional epistemic overlay handles uncertainty over order. A further extension of the model would incorporate a process matrix for the order of measurement, which need not be fixed and could allow an indefinite causal structure (Oreshkov, Costa, and Brukner, 2012; Chiribella et al., 2013). We hope the results reported in this paper are a first step towards these important extensions.
Another open direction is to allow Alice and Bob to undertake measurements described by POVMs in place of the PVMs assumed in this paper. Now, the post-measurement state can depend not only on the outcome but also on the particular quantum instrument realizing the POVM (Gühne et al., 2023). Information about these instruments might or might not be accessible to the agents, and, in another future extension, we would need to enrich the epistemics further to account for this feature.
@@
Figure 2 shows that common certainty of probability assessments need not coincide with the shared assessment Alice and Bob would obtain after pooling their information. Here, Alice’s (resp. Bob’s) partition is the rows (resp. columns). The event of interest $E$ is $\{\omega_{1},\omega_{4}\}$. At any state, it is common certainty that Alice and Bob both assign probability $1/2$ to $E$. Yet, if they pooled their information, their shared assessment would be $1$ or 0, depending on the true state.
@@
The next lemma ensures that all conditional probabilities in the common-certainty recursion are well defined (there are no vanishing denominators).
@@
Since $C_{*}=A_{*}B_{*}$ is the projector onto $\operatorname{ran}(A_{*})\cap\operatorname{ran}(B_{*})$, we have $0\leq C_{*}\leq A_{*}\leq A_{n}$ in the Loewner (positive semidefinite) order, and similarly $0\leq C_{*}\leq B_{*}\leq B_{n}$. Monotonicity of the trace with respect to the operator order and $\rho\geq 0$ give the inequalities. Strict positivity is guaranteed by Condition 4 in the main text. ∎